OpenStudy (anonymous):

an orchard has 2030 trees. The number of trees in each row is 12 more than twice the number of rows. How many trees are in each row?

Directrix (directrix):

Let R = the number of rows Let T = the number of trees per row T = 12 + 2R because of this: The number of trees in each row is 12 more than twice the number of rows. R*T = 2030

Directrix (directrix):

So, solve these two equations simultaneously by the substitution method: T = 12 + 2R R*T = 2030 ------------- R * ( 12 + 2R) = 2030 Solve for R. @melissalovee